Enter 1000 lb / h of oil at 100 ° F in a 1-inch copper tube, the length of the tube is 12 ft. If the tube surface is maintained at 220 ° F, determine the oil outlet temperature.

Cp=0.5 Btu/lbm °F p=55 lbm/ft M = 1.5 lbm/ft.h k=0.10 Btu/h. ft. °F
Answer #1

22222222222222 mass m= 1000 lb/h = 0.277 lbs Ti = 100°F D = 1 in L = 12 tt Tw=220°f. Re: PVD M = m = PAV m = p 77 D2 V 4 m =using pittus - Boelter equation with oll belung Nud = 0.023 (Red)** (po 30.4 heated Nud = 0.023 (10185.9163) (75):4 Nud = 82.1.86416 = 120 220-te 220-te = 1,8616 220- Te - 64.268 Te = 155.73F